### v1

@@ 3.2.2 Photometric Uncertainty
Figure 9 demonstrates the photometric uncertainty of the selected member stars in each band. While the shallowest -band photometry show relatively large uncertainty, the other three bands achieve 1% level precision roughly up to 21.5, 22, 21.5 mag in the , , bands, respectively. The multiple quantized sequences seen in each band reflect variation in image depth across the field. Photometry from deeper exposures (i.e., more input images) yields smaller uncertainties. To illustrate this, each panel is color-coded by the number of input images contributing to the measurement at each object’s position. The distinct sequences correspond to different numbers of input images, confirming that depth is the dominant factor shaping the structure of these sequences.
@@ 3.2.3 Analysis of Photometric Scatter
The simple stellar population of 47 Tuc affords the opportunity to diagnose possible sources of scatter in the DP1 photometry. Figure 10 shows the vs. color-color diagram for bright stars (), which features a tight linear sequence. We considered four possible sources to explain the scatter about this sequence: 1) PSF-fitting errors stemming from known Poisson and other observational noise sources; 2) crowding errors (e.g., Olsen et al., 2003); 3) differential extinction by dust (e.g., Pancino et al., 2024); and 4) unknown sources of observational systematic error.
@@
We do not claim to have measured differential extinction in 47 Tuc based on modeling the photometric scatter in DP1, but do conclude that in the areas covered, differential extinction can not be as large as seen in Pancino et al. (2024). Those authors cited camera-specific systematic errors as limiting their measurement, and we cannot rule out such errors in DP1. Figure 13 shows the observed median values of the offsets about the line fit to the vs. plane. The figure shows spatially coherent variations in at scales of deg, which matches the scale of the patches in the DP1 reduction. There is not, however, obvious correlation with patch boundaries; more data will be needed to understand the source of photometric scatter in 47 Tuc.
@@ 3.3 Variables in the 47 Tuc Field
As an assessment of the pipeline performance and DP1 data products, we crossmatched the 47 Tuc Object and DiaObject (NSF-DOE Vera C. Rubin Observatory, 2025e) tables to a catalog of known variables from Weldrake et al. (2004). The original catalog contains 100 variables stars across an approximately 1 square degree field centered on the cluster, which is presented in their paper in Table 2. Their Tables 3–5 then detail the fitted period and amplitude of of each variable, broken out by variable type: RR Lyrae (RRLyr), Eclipsing Binaries (EcBs) and others, mostly long period variables (LPVs). Separately, a table was published to SIMBAD containing additional photometry, as well as an identifier column and links to other references for each variable. We consolidated the tables, ignoring rows with only limits or approximations of the period, amplitude or photometry, leaving 72 variables remaining. The 72 variables left include 19 EcBs, 8 detached EcBs, 11 LPVs, and 36 RRLyr. All RRLyr recovered likely belong to the SMC according to distance moduli derived by Weldrake et al. (2004).
@@
We then used the Large Survey Database (Caplar et al., 2025, LSDB) to crossmatch the 72 variable catalog to both the Object table and DiaObject table narrowed down to contain only those located in the 47 Tuc field. We find 5 matches to Objects and 62 matches with DiaObjects. Three variables, all RR Lyrae (ID: V23, V73, and V88) are found in both tables. The two objects not found in the DiaObject table correspond to one EcB and one detached EcB. As discussed in Section 2.2.2, the skipped debelending on the denser areas closer to the cluster core results in fewer detected objects toward the center, as shown in the middle panel of Figure 2. In contrast, more DiaObjects are detected closer to the cluster core because the difference images are less crowded after template subtraction. For our variable star crossmatching, we therefore use the DiaObject table, particularly because we expect there to be residual flux in the difference images due to the variability of our targets.
@@
In Figure 14, we show the science, template, and difference images in a small field within 47 Tuc from visit 2024112400101 and detector 3. The overplotted cyan crosses show the diaSources detected in the difference image. The source marked with the red circle is a known RR Lyrae variable star and shows a clear negative flux in the difference image. It appears some of the diaSources may be incomplete subtractions of the brightest stars. At least one of the detections appears to be on a diffraction spike of the brightest star in the image.
@@
There are several flux measurements available in the DP1 datasets. The scienceFlux from the DiaSource table is the forced PSF flux measured on the direct image at the position of the diaSource. We also have forced photometry measurements at the position of the associated diaObject on the difference image (psfDiffFlux) and the preliminary visit image (psfFlux) available from the ForcedSourceOnDiaObject table.
@@
In Figure 15, we plot one of the RRLyr stars (diaObjectId = 579584602276888841; ID: V23) recovered in our crossmatch. Because the 47 Tuc field was only visited on 4 days (Table 1), the lightcurve is quite sparse. However, the shape of the lightcurve looks like a typical RR Lyrae profile and matches the phase-wrapped lightcurve of the same source from Weldrake et al. (2004). We plotted the forced photometry on the visit images at the diaObject position (psfFlux), as there are more flux measurements in the ForcedSourceOnDiaObject table than in the DiaSource table. This is because the source is not always detected in the difference images, for example when its science flux is close to the flux in the template image, and thus does not always make it into the DiaSource table. The phase-folded lightcurve shows a consistent pattern with the densely-sampled one from Gaia DR3 photometry (Clementini et al., 2023).
@@ 4 Summary
This study presents a detailed exploration and constructive assessment of Rubin Observatory DP1 data centered on the globular cluster 47 Tuc. The LSSTComCam imaging provided valuable early photometric measurements, while also revealing challenges from crowding, particularly near the core of 47 Tuc and toward the SMC. Although crowding affected deblending and subsequent measurements, improvements to scarlet lite are anticipated to enhance processing capabilities in crowded regions before the first data release.
@@
Further, we successfully crossmatched known variable stars within the 47 Tuc field against the DP1 data, recovering three RR Lyrae stars and two eclipsing binaries. Despite sparse temporal sampling, forced photometry at their fixed locations in every image aligns pretty well with the denser lightcurves from the literature, underscoring Rubin’s promising capabilities for variability detections even in crowded regions.
Overall, while challenges remain, the DP1 data around 47 Tuc convincingly showcase Rubin Observatory’s strong potential for detailed stellar population and variability analyses in crowded stellar fields. Continued improvements to the Rubin Science Pipelines and in-kind programs dedicated to crowded-field stellar photometry are expected to deliver even higher-quality results in future DP2 and DR1.

### v2

@@ 3.2.2 Photometric Uncertainty
Figure 9 demonstrates the photometric uncertainty of the selected member stars in each band. While the shallowest -band photometry show relatively large uncertainty, the other three bands achieve 1% level precision roughly down to 21.5, 22, 21.5 mag in the , , bands, respectively. The multiple quantized sequences seen in each band reflect variation in image depth across the field. Photometry from deeper exposures (i.e., more input images) yields smaller uncertainties. To illustrate this, each panel is color-coded by the number of input images contributing to the measurement at each object’s position. The distinct sequences correspond to different numbers of input images, confirming that depth is the dominant factor shaping the structure of these sequences.
@@ 3.2.3 Analysis of Photometric Scatter
The “simple” stellar population of 47 Tuc affords the opportunity to diagnose possible sources of scatter in the DP1 photometry. Figure 10 shows the vs. color-color diagram for bright stars (), which features a tight linear sequence. We considered five possible sources to explain the scatter about this sequence: 1) PSF-fitting errors stemming from known Poisson and other observational noise sources; 2) crowding errors (e.g., Olsen et al., 2003); 3) differential extinction by dust (e.g., Pancino et al., 2024); 4) multiple stellar populations, as observed by Milone et al. (2012) and Marino et al. (2023); and 5) unknown sources of observational systematic error.
@@
We do not claim to have measured differential extinction in 47 Tuc based on modeling the photometric scatter in DP1, but we do conclude that in the areas covered, differential extinction can not be as large as seen in Pancino et al. (2024). Those authors cited camera-specific systematic errors as limiting their measurement, and we cannot rule out such errors in DP1. Figure 13 shows the observed median values of the offsets about the line fit to the vs. plane. The figure shows spatially coherent variations in at scales of deg, which matches the scale of the patches in the DP1 reduction. There is not, however, obvious correlation with patch boundaries; more data will be needed to understand the source of photometric scatter in 47 Tuc.
Multiple stellar populations are also a candidate to explain the scatter that we see. They are known to have radial dependence (e.g., Milone et al., 2012), but would not be expected to show the azimuthal variations that we see in Figure 13. With more data from LSSTCam observations of 47 Tuc, we should have an excellent opportunity to understand the source of photometric scatter in 47 Tuc, including the presence of differential extinction and multiple stellar populations.
@@ 3.3 Variables in the 47 Tuc Field
As an assessment of the pipeline performance and DP1 data products, we crossmatched the 47 Tuc field’s Object and DiaObject (NSF-DOE Vera C. Rubin Observatory, 2025e) tables to a catalog of known variables from Weldrake et al. (2004). The original catalog contains 100 variables stars across an approximately 1 deg2 field centered on the cluster, which is presented in their paper in Table 2. Their Tables 3–5 then detail the fitted period and amplitude of of each variable, broken out by variable type: RR Lyrae (RRLyr), Eclipsing Binaries (EcBs) and others, mostly long period variables (LPVs). Separately, a table was published to SIMBAD containing additional photometry, as well as an identifier column and links to other references for each variable. We consolidated the tables, ignoring rows with only limits or approximations of the period, amplitude or photometry, leaving 72 variables remaining. The 72 variables left include 19 EcBs, 8 detached EcBs, 11 LPVs, and 36 RRLyr. All RRLyr recovered likely belong to the SMC according to distance moduli derived by Weldrake et al. (2004).
@@
We then used the Large Survey Database (LSDB; Caplar et al., 2025) to crossmatch the 72 variable catalog to both the Object table and DiaObject table narrowed down to contain only those located in the 47 Tuc field. We find 5 matches to the Object table and 62 matches with the DiaObject table. Three variables, all RR Lyrae (ID: V23, V73, and V88), are found in both tables. The two objects not found in the DiaObject table correspond to one EcB and one detached EcB. As discussed in Section 2.2.2, the skipped debelending on the denser areas closer to the cluster core results in fewer detected objects toward the center, as shown in the middle panel of Figure 2. In contrast, more “DiaObjects” are detected closer to the cluster core because the difference images are less crowded after template subtraction. For our variable star crossmatching, we therefore use the DiaObject table, particularly because we expect there to be residual flux in the difference images due to the variability of our targets.
@@
In Figure 14, we show the science, template, and difference images in a small field within 47 Tuc from visit 2024112400101 and detector 3. The overplotted cyan crosses show the “DiaSources” detected in the difference image. The source marked with the red circle is a known RR Lyrae variable star and shows a clear negative flux in the difference image. It appears some of the DiaSources may be incomplete subtractions of the brightest stars. At least one of the detections appears to be on a diffraction spike of the brightest star in the image.
@@
There are several flux measurements available in the DP1 datasets. The scienceFlux column from the DiaSource table is the forced PSF flux measured at the position of the DiaSource in all visit images where the DiaSource is detected. We also have forced photometry measurements at the position of the associated DiaObject on the difference image (psfDiffFlux column) and the visit image (psfFlux column) available from the ForcedSourceOnDiaObject table.
@@
In Figure 15, we plot one of the RRLyr stars (diaObjectId = 579584602276888841; ID: V23) recovered in our crossmatch. Because the 47 Tuc field was only visited on 4 days (Table 1), the lightcurve is quite sparse. However, the shape of the lightcurve looks like an RR Lyrae profile and matches the phase-wrapped lightcurve of the same source from Weldrake et al. (2004). We plotted the forced photometry on the visit images at the DiaObject position (psfFlux), as there are more flux measurements in the ForcedSourceOnDiaObject table than in the DiaSource table. This is because the source is not always detected in the difference images, for example when its science flux is close to the flux in the template image, and thus does not always make it into the DiaSource table. The phase-folded lightcurve shows a consistent pattern with the densely-sampled one from Gaia DR3 photometry (Clementini et al., 2023).
@@ 4 Summary
This study presents a detailed exploration and constructive assessment of Rubin Observatory DP1 data centered on the globular cluster 47 Tuc. The LSSTComCam imaging provided valuable early photometric measurements, while also revealing challenges from crowding, particularly the inner region of 47 Tuc and toward the SMC. Although crowding affected deblending and subsequent measurements, improvements to scarlet lite are anticipated to enhance processing capabilities in crowded regions before the first data release.
@@
Further, we successfully crossmatched known variable stars within the 47 Tuc field against the DP1 data, recovering three RR Lyrae stars and two eclipsing binaries in the coadd-based object catalog, and identify 62 of the 72 known variables in the difference image-based object catalog. Despite sparse temporal sampling, forced photometry at their fixed locations in every image aligns pretty well with the denser lightcurves from the literature, underscoring Rubin’s promising capabilities for variability detections even in crowded regions.
Overall, while challenges remain, the DP1 data around 47 Tuc convincingly showcase Rubin Observatory’s strong potential for detailed stellar population and variability analyses in crowded stellar fields. Continued improvements to the Rubin Science Pipelines and in-kind programs dedicated to crowded-field stellar photometry are expected to deliver even higher-quality results in future DP2 and DR1. The substantial increase in depth and cadence over the 10-yr LSST survey will enable more complete variability studies and high-fidelity characterization of stellar populations in 47 Tuc and similar dense fields.